Multiple Products, Break-Even Analysis, Operating Leverage

Carlyle Lighting Products produces two different types of lamps: a floor lamp and a desk lamp. Floor lamps sell for $30, and desk lamps sell for $20. The projected income statement for the coming year follows:

Sales $600,000
Total variable cost 400,000
Contribution margin $200,000
Total fixed cost 150,000
Operating income $50,000

The owner of Carlyle estimates that 60% of the sales revenues will be produced by floor lamps and the remaining 40% by desk lamps. Floor lamps are also responsible for 60% of the variable cost. Of the fixed cost, one-third is common to both products, and one-half is directly traceable to the floor lamp product line.

Required:

1. Compute the sales revenue that must be earned for Carlyle to break even. Round the contribution margin ratio to six decimals and sales revenue to the nearest dollar.
$

2. Compute the number of floor lamps and desk lamps that must be sold for Carlyle to break even. Round variable rates and contribution margins to four decimal places in your calculations. Round the final answers to the nearest whole dollar.

Floor lamps units
Desk lamps units

3. Compute the degree of operating leverage for Carlyle.

Now assume that the actual revenues will be 40% higher than the projected revenues. By what percentage will profits increase with this change in sales volume?
%
Answer #1

1. Break even = fixed cost / contribution margin

= 150,000 / 33.33% = $ 450,000

Explanation:

PV Ratio = Contribution / Sales * 100

= 200,000 / 600,000 * 100 = 33.33%

2.Total basket contribution margin = $ 16.67

Break-even number of floor lamps = 9,000

Break-even number of desk lamps = 9,000

Explanation:

Floor Lamp Desk Lamp

Selling Price $ 30.00 $ 20.00

Less: Variable Cost $ 20.00 $ 13.33

Contribution $ 10.00 $ 6.67

Sales mix 50%. 50%

Weighted Contribution margin $ 8.33

Fixed cost $ 150,000

Break-even point 18,000

Sale from each product 9,000 9,000

3.

Operating Leverage = Contribution / Operating Income

= 200,000 / 50,000 = 4

Change in profits = 40% * 4 = 160%